# Some new findings concerning value distribution of a pair of delay-differential polynomials

### v3

@@ 2. Hayman’s conjecture on a pair of delay-differential polynomials
Let $f$ and $g$ be both transcendental entire functions, $\min\{n,m\}\geq 2$, $\alpha$ be a non-zero small function with respect to $f$ and $g$. Then, $f^{n}L(z,g)$ and $g^{m}L(z,f)$ cannot have a common non-zero generalized Picard exceptional small function $\alpha$, where $L(z,f)$ and $L(z,g)$ are non-zero linear differential polynomial with small functions of $f$ and $g$ as its coefficients.
@@
If $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,f)$ and $L(z,g)$ are non-zero linear delay-differential polynomial with small functions of $f$ and $g$ as their coefficients, $\alpha$ is a non-zero small funtion with respect to $f$ and $g$, then $f^{n}L(z,g)$ and $g^{n}L(z,f)$ cannot have a common non-zero generalized Picard exceptional small function $\alpha$ by using the similar proof of Theorem 2.1. Combining this result with Theorem 2.1, we give a positive answer to Question 1 for entire functions $f$ and $g$.
@@
    If $\min\{m,n\}\geq 2$, then $f^{n}L(z,g)$ and $g^{m}(z)L(z,f)$ cannot have simultaneously non-zero generalized Picard exceptional small functions, where $L(z,f)$ and $L(z,g)$ are non-zero linear differential polynomial with small functions of $f$ and $g$ as its coefficients.
@@
If $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,f)$ and $L(z,g)$ are linear delay-differential polynomial with small functions of $f$ and $g$ as its coefficients, then Theorem 2.2-$(1)$ still holds by using the similar proof of Theorem 2.2.
@@
We only need to consider $m(r,g)\leq m(r,f)$, since the case $m(r,f)\leq m(r,g)$ can be studied by the same methods. Below, we will consider two subcases: $m(r,g)\leq m(r,f)$ and whether $\alpha$ is a small function of $f_{2}$ or not.
@@
$m(r,g)\leq m(r,f)$ and $\alpha$ is a small function of $f_{2}$. Since $m(r,g)\leq m(r,f)$, then from (2), we get $(n+1)m(r,f)+S(r,f)\geq m(r,f_{1})\geq(n-1)m(r,f)$, which means $\alpha$ is a small function of $f_{1}$. We claim at least one of $f^{n}L(z,g)-\alpha$ and $g^{m}L(z,f)-\alpha$ has infinitely many zeros. Otherwise, suppose $f^{n}L(z,g)-\alpha$ and $g^{m}L(z,f)-\alpha$ have finitely many zeros. Since $\alpha$ is a small function of $f_{1}$ and $f_{2}$, by the Weierstrass’s factorization theorem, we have
@@
$m(r,g)\leq m(r,f)$ and $\alpha$ is not a small function of $f_{2}$. We claim $f^{n}L(z,g)-\alpha$ has infinitely many zeros. Otherwise, $f^{n}L(z,g)-\alpha$ has finitely many zeros. As the same as the Case 1, we get equations (3.1)-(2) and $f(z)=A_{3}(z)e^{B_{3}(z)}$, where $A_{3}$ and $B_{3}$ are entire function, such that $T(r,A_{3})=S(r,f)$. Since $\alpha$ is not a small function of $f_{2}$, then $T(r,f_{2})\leq O(T(r,\alpha))=S(r,f)$.
By applying the logarithmic derivative lemma, we can obtain $m(r,\frac{L(z,f)}{f})\leq S(r,f)$. Since $N(r,\frac{1}{f})=S(r,f)$, we have $T(r,\frac{L(z,f)}{f})=m(r,\frac{L(z,f)}{f})+N(r,\frac{L(z,f)}{f})=S(r,f)$. Thus $\frac{L(z,f)}{f}=a_{5}$, where $a_{5}$ is a small function of $f$. Therefore $g^{m}=\frac{f_{1}}{L(z,f)}=f_{1}\frac{1}{a_{5}f}$, which means $g=A_{4}e^{\frac{-B_{3}}{m}}$, which $A_{4}$ is a small function of $f$. As the same as $L(z,f)$, we have $L(z,g)=a_{6}g$, where $a_{6}$ is a small function of $g$.
@@
4. Discussion
It should be noted that all the results presented in this article are still valid when $f$ and $g$ are meromorphic functions and satisfy the conditions $N(r,f)=S(r,f)$ and $N(r,g)=S(r,g)$. However, these results do not necessarily apply to general meromorphic functions $f$ and $g$. For example, in the proof of Theorem 2.1, if $N(r,f)\not=S(r,f)$, then $p(z)$ in (3.1) is not a small function of $f$. Consequently, we cannot find a zero $z_{0}$ of $f$ such that the coefficients of the equation (3.2) are neither zero nor infinite at that point.

### v5

@@ 2. Hayman’s conjecture on a pair of delay-differential polynomials
Let $f$ and $g$ be both transcendental entire functions, $\min\{n,m\}\geq 2$, $\alpha$ be a non-zero small function with respect to $f$ and $g$. Then, $f^{n}L(z,g)$ and $g^{m}L(z,f)$ cannot have a common non-zero generalized Picard exceptional small function $\alpha$, where $L(z,f)$ and $L(z,g)$ are non-zero linear differential polynomial with small entire functions of $f$ and $g$ as its coefficients.
@@
If $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,f)$ and $L(z,g)$ are non-zero linear delay-differential polynomial with small entire functions of $f$ and $g$ as their coefficients, $\alpha$ is a non-zero small funtion with respect to $f$ and $g$, then $f^{n}L(z,g)$ and $g^{n}L(z,f)$ cannot have a common non-zero generalized Picard exceptional small function $\alpha$ by using the similar proof of Theorem 2.1. Combining this result with Theorem 2.1, we give a positive answer to Question 1 for entire functions $f$ and $g$.
@@
    If $\min\{m,n\}\geq 2$, then $f^{n}L(z,g)$ and $g^{m}(z)L(z,f)$ cannot have simultaneously non-zero generalized Picard exceptional small functions, where $L(z,f)$ and $L(z,g)$ are non-zero linear differential polynomial with small entire functions of $f$ and $g$ as its coefficients.
@@
If $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,f)$ and $L(z,g)$ are linear delay-differential polynomial with small entire functions of $f$ and $g$ as its coefficients, then Theorem 2.2-$(1)$ still holds by using the similar proof of Theorem 2.2.
@@
We only need to consider $m(r,g)\leq m(r,f)$, since the case $m(r,f)\leq m(r,g)$ can be studied by the same methods.
Since $m(r,g)\leq m(r,f)$, then from (2), we get $(n+1)m(r,f)+S(r,f)\geq m(r,f_{1})\geq(n-1)m(r,f)$, which means $\alpha$ is a small function of $f_{1}$.
Below, we will consider two subcases: $m(r,g)\leq m(r,f)$ and whether $f_{2}$ is a small function of $f$ or not.
@@
$m(r,g)\leq m(r,f)$ and $f_{2}$ is not a small function of $f$. Since $f_{2}$ is not a small function of $f$, then $T(r,f_{2})=O(T(r,f))$, $\alpha$ is a small function of $f_{2}$. We claim at least one of $f^{n}L(z,g)-\alpha$ and $g^{m}L(z,f)-\alpha$ has infinitely many zeros. Otherwise, suppose $f^{n}L(z,g)-\alpha$ and $g^{m}L(z,f)-\alpha$ have finitely many zeros. Since $\alpha$ is a small function of $f_{1}$ and $f_{2}$, by the Weierstrass’s factorization theorem, we have
@@
$m(r,g)\leq m(r,f)$ and $f_{2}$ is a small function of $f$. We claim $f^{n}L(z,g)-\alpha$ has infinitely many zeros. Otherwise, $f^{n}L(z,g)-\alpha$ has finitely many zeros. As the same as the Case 1, we get equations (3.1)-(2) and $f(z)=A_{3}(z)e^{B_{3}(z)}$, where $A_{3}$ and $B_{3}$ are entire function, such that $T(r,A_{3})=S(r,f)$.
By applying the logarithmic derivative lemma, we can obtain $m(r,\frac{L(z,f)}{f})\leq S(r,f)$. Since $N(r,\frac{1}{f})=S(r,f)$, we have $T(r,\frac{L(z,f)}{f})=m(r,\frac{L(z,f)}{f})+N(r,\frac{L(z,f)}{f})=S(r,f)$. Thus $\frac{L(z,f)}{f}=a_{5}$, where $a_{5}$ is a small function of $f$. Therefore $g^{m}=\frac{f_{2}}{L(z,f)}=f_{2}\frac{1}{a_{5}f}$, which means $g=A_{4}e^{\frac{-B_{3}}{m}}$, which $A_{4}$ is a small function of $f$. As the same as $L(z,f)$, we have $L(z,g)=a_{6}g$, where $a_{6}$ is a small function of $g$.